# Diffusion in driven periodic systems: from weak to strong damping regimes

### v1

@@ III.2 Diffusion versus friction: non-monotonic dependence
For $\gamma>\gamma_{2}$ we observe a decrease in the spatial interval $\Delta x$ of oscillations around the minimum. In other words, oscillations are focused closer and closer to the minimum of the potential as $\gamma$ grows. As a result, the effective potential barrier height increases and the diffusion coefficient decreases. To sum it up, the sequence of panels in Fig. 5 reveals the mechanism which is responsible for the non-monotonic behavior of diffusion coefficient $D$ in the strong damping regime.
IV Vibrational mechanics approach
The diffusional behavior for strong damping can be satisfactory explained in terms of the effective potential barrier for the particle motion in a spatially periodic potential. To this aim, we use vibrational mechanics methods and replace the starting Eq. (3) by the following Langevin equation boroEPL ; march-prl
@@ IV Vibrational mechanics approach
where $\tilde{x}=\tilde{x}(t)$ is a slow part of the dynamics (the fast one is represented by free oscillations) and $J_{0}(\psi)$ is the zero order Bessel function specialF of the argument
$\psi=\frac{a}{\omega\sqrt{\omega^{2}+\gamma^{2}}}.$ (13)
We note that the starting total time-dependent potential $V(x,t)=-\cos(x)-ax\sin(\omega t+\phi_{0})$ is replaced by the time-independent renormalized potential $V_{eff}(x)=-J_{0}(\psi)\cos(x)$. In this approach, the particle diffuses in a renormalized potential of the effective barrier $2J_{0}(\psi)$. This approximation is postulated to be correct for the high frequency limit $\omega\gg 1$. In this section, we want to verify to what extend this approximation reconstructs the exact dynamics. In Fig. 6 we compare results obtained from the exact dynamics (3) and its approximate counterpart (12).
For the case of weak dissipation, we can apply the modified equation (9) which reads pavli ; entropy23
@@
where $d_{0}=J_{0}(\psi)$. In panel (b) of Fig. 6 we compare the dependence $D(\gamma)$ from this relation with the exact results of simulations. The discrepancy is very big and (14) cannot reconstruct the local maximum for $a=14$ which is at $\gamma=0.06$. Therefore this approximation totally fails.
However, for strong damping, the approximate expression for the diffusion coefficient can be obtained from the modified equation (10) which takes the following form lifson ; entropy23
$D=\frac{Q}{\gamma I_{0}^{2}(J_{0}(\psi)/Q)}.$ (15)
Figure 6: Diffusion $D$ versus the friction coefficient $\gamma$ depicted for the 1st (panel (a)) and the 2nd (panel (b)) maximum and minimum of $D$ taken from Fig. 1 together with a comparison of the numerical results and the approximate formula given by Eq. (14) and (15). The ac-driving frequency reads $\omega=1.59$ while temperature is $Q=0.05$.
In this case, the approximation (15) correctly reproduces the exact results for the friction coefficient $\gamma>0.3$ which could be treated as the regime of strong damping. Moreover, it turns out that for larger values of the amplitude $a$ new maxima emerge, e.g. for $a=18$ there are two maxima at $\gamma_{M}=1.05$ and $\gamma_{M}=1.4$ but for $a=35$ four maxima exist at $\gamma_{M}=0.95,2,3.75,9.5$ (not shown). The envelope of these points behaves roughly like $D(\gamma_{M})\sim 1/\gamma_{M}$.
V Summary
@@
We showed that the dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ characterizing the dissipation mechanism can render unexpected features. Typically, diffusion $D$ is a decreasing function of the damping $\gamma$. In contrast, we demonstrated that for an archetypal model of nonequilibrium system consisting of a Brownian particle moving in a symmetric and spatially periodic potential subjected to an unbiased time-periodic driving the diffusion coefficient $D$ can grow when the damping $\gamma$ increases. Moreover, for sufficiently large amplitude of the ac-driving the characteristics $D(\gamma)$ can exhibit several local maxima and minima in the strong damping regime.
We revealed the mechanisms standing behind this counterintuitive phenomenon. In the weak damping limit the maximum of diffusion is induced by a non-monotonic change of population of the running states when the dissipation increases. In contrast, in the strong friction regime it is related to vanishing of the effective barrier of the potential in which the Brownian particle dwells. The latter effect is caused by displacement of the location (center) and amplitude of the locked orbits induced by the damping change. Properties in the strong dissipation regime are well reproduced by the vibrational mechanics methods proposed in Refs. boroEPL ; march-prl .
Acknowledgments
Appendix A Nondimensionalization of the Langevin equation
@@
The starting dimensional Langevin equation has the form entropy23 ; marchPRE23 ; spiechowicz2016njp
@@
$F(t)=A\sin{(\Omega t)}.$ (18)
Thermal equilibrium fluctuations describing the influence of thermostat of temperature $T$ are normally assumed as $\delta$-correlated Gaussian white noise of zero-mean value,
@@
The characteristic length $L$ is the period of the potential $U(x)$ and the characteristic time $\tau_{0}$ is related to the period of small oscillations inside the potential well of $U(x)$.
@@
$\ddot{\hat{x}}+\gamma\dot{\hat{x}}=-\sin{\hat{x}}+a\sin(\omega\hat{t})+\sqrt{2\gamma Q}\,\hat{\xi}(\hat{t}),$ (21)
@@
$\gamma=\frac{\Gamma}{M}\tau_{0},\quad a=\frac{1}{2\pi}\frac{L}{\Delta U}A,\quad\omega=\tau_{0}\Omega,\quad Q=\frac{k_{B}T}{\Delta U}.$
The dimensionless potential $V(x)$ of the period $L=2\pi$ is $V(\hat{x})=U((L/2\pi)\hat{x})/\Delta U=-\cos{\hat{x}}$ and the corresponding potential force reads $f(x)=-\hat{U}^{\prime}(\hat{x})=-\sin\hat{x}$. The rescaled thermal noise is $\hat{\xi}(\hat{t})=(L/2\pi\Delta U)\xi(t)=(L/2\pi\Delta U)\xi(\tau_{0}\hat{t})$ and has the same statistical properties as $\xi(t)$; i.e., $\langle\hat{\xi}(\hat{t})\rangle=0$ and $\langle\hat{\xi}(\hat{t})\hat{\xi}(\hat{s})\rangle=\delta(\hat{t}-\hat{s})$. The dimensionless noise intensity $Q$ is the ratio of thermal energy and half of the activation energy the particle needs to overcome the nonrescaled potential barrier. In order to simplify the notation we omit the hat-notation in Eq. (3).

### v2

@@ A. Giant enhancement of diffusion
For $\gamma>\gamma_{2}$ we observe a decrease in the spatial interval $\Delta x$ of oscillations around the minimum. In other words, oscillations are focused closer and closer to the minimum of the potential as $\gamma$ grows. As a result, the effective potential barrier height increases and the diffusion coefficient decreases. To sum it up, the sequence of panels in Fig. 3 reveals the mechanism which is responsible for the non-monotonic behavior of diffusion coefficient $D$ in the strong damping regime. We should be aware of limitations of this picture because in the noisy system its dynamics is much more complicated.
In panel (g) of Fig. 3 we collect the above information and extend it to continuous change of $\gamma$. It has the form of the bifurcation diagram for the centers of periodic orbits $x_{c}$ as a function of the friction coefficient $\gamma$. The crucial point is emergence of two classes of periodic orbits for $\gamma>3.2$. For the first class the centers $x_{c}$ move to the minimum $x=0$ of the potential $W(x)$ whereas for the second $x_{c}$ goes to its another minimum $x=2\pi$. It is visualized as a pitchfork bifurcation and reveals the mechanism which is responsible for the non-monotonic behavior of diffusion coefficient $D$ versus dissipation in the strong damping regime.
Figure 4: The weak dissipation regime: Panel (a) $D$ versus $\gamma$. Inset: The probability distribution $P(V)$ of the period-averaged velocity $V$ of the Brownian particle. Panel (b) The stationary probability P of the Brownian particle to reside in various states $V=0,V=\pm k\omega\}$ characterizing the deterministic counterpart of the studied setup. The black line corresponds to the sum of probabilities for all running states. Parameters are: $a=14$, $\omega=1.59$ and $Q=0.05$.
B. Diffusion in weak dissipation regime
The non-monotonic dependence of diffusion on friction is also detected in the weak damping regime. In Fig. 4 we depict this behavior. In the vicinity of $\gamma=0.06$, the probability distribution $P(V)$ for the period-averaged velocity $V$ of the Brownian particle has three remarkable maxima, one corresponding to the locked state $V=0$ and two symmetric running states of the averaged velocities $V=\pm\omega$. At non-zero temperature, there are random transitions induced by thermal fluctuations between these states. For one trajectory, the particle can follow the running solution, next for some time interval it resides in the localized state and later it again travels along the running state either in the same or the opposite direction. In this way, for a given trajectory various deterministic solutions are visited in a random sequence. In consequence the deterministic structure of locked and running states impacts the spread of trajectories.
In panel (b) of Fig. 4 we present the probability P to reside in various states $\{V=0,\pm k\omega\}$ characterizing the deterministic counterpart of the studied setup. It allows us to answer the question about the mechanism standing behind the observed pronounced maximum of the diffusion coefficient $D$ as a function of the friction $\gamma$ for weak damping regime $\gamma\approx 0.06$. The initial decay of this characteristics is clearly related to a decrease of the running states $V=\pm k\omega$ and simultaneous grow of the locked ones $V=0$. Next, there is one-to-one correspondence between the non-monotonic behavior of $D$ and the probability P to reside in the running states. Therefore the local growth of the number of running trajectories is responsible for the increase of $D$ despite the fact that the damping $\gamma$ is getting bigger as well. We note that this mechanism of growing of $D$ is radically different than in the case of the strong dissipation regime.
In Fig. 5, we put together results of Fig. 2 and Fig. 4 and depict a spread of trajectories for three values of the friction $\gamma$. In panel (b), it is the weak dissipation regime with $\gamma=0.06$ (the first maximum of $D$) in which both running and locked trajectories emerge. In panel (c) $\gamma=0.9$ and only locked trajectories exist. In panel (d), the strong dissipation regime with $\gamma=3.3$ is presented for which there are only locked trajectories likewise. For this value of $\gamma$ the second maximum of $D$ is detected. We do not present a corresponding figure at the minimum $\gamma=1.66$ of the diffusion coefficient $D$ due to the very small spread of trajectories.
C. Vibrational mechanics approach
It is argued that within vibrational mechanics methods sorokin the high-frequency dynamics (3) can be approximated by a much simpler Langevin equation. Following the approach of boroEPL ; march-prl we can separate
$x(t)={\tilde{x}}(t)+\psi\sin(\omega t+\phi_{1}),$ (8)
where $\tilde{x}(t)$ represents a slowly time-modulated stochastic process. For $\omega\gg 1$, the fast oscillating terms can be averaged over the period of the ac driving and the Langevin equation (3) for the slow reduced spatial variable $\tilde{x}(t)\equiv\tilde{x}$ can be written as
@@ C. Vibrational mechanics approach
where $J_{0}(\psi)$ is the zero order Bessel function specialF of the argument
$\psi=\psi(\gamma)=\frac{a}{\omega\sqrt{\omega^{2}+\gamma^{2}}}.$ (10)
We now want to apply Eq. (9) to test its correctness in both strong and weak damping regimes for the case consider above, i.e. for the frequency $\omega=1.59$, which obviously does not fulfill the condition $\omega\gg 1$. We note that the starting total time-dependent potential $\mathbb{W}(x,t)=-\cos(x)-ax\sin(\omega t+\phi_{0})$ is replaced by the time-independent renormalized potential $\mathbb{W}_{\mbox{eff}}(x)=-J_{0}(\psi)\cos(x)$. In this approach, the particle diffuses in a renormalized potential of the effective barrier $2J_{0}(\psi)$. In Fig. 5 we compare results obtained from the exact dynamics (3) and its approximate counterpart (9).
Figure 5: The diffusion coefficient $D$ versus the friction coefficient $\gamma$. Comparison of the exact numerical and approximate results given by Eqs. (11) (strong damping: dotted line) and (12) (weak damping: dotted-dashed line). The parameters are: $a=14$, $\omega=1.59$ and $Q=0.05$. Exemplary set of the system trajectories vs. time $t/\mathsf{T}$ ($\mathsf{T}=2\pi/\omega$) for $\gamma=0.06$, the first maximum (panel (b); $\gamma=1$, (panel c) and $\gamma=3.3$, the second maximum (panel d). Note that the vertical axes have different scales in the panels (b)-(d).
For strong damping, the approximate expression for the diffusion coefficient can be obtained from Eqs. (26) and (27), which takes now the following form lifson ; entropy23
$D=\frac{Q}{\gamma I_{0}^{2}(J_{0}(\psi)/Q)},$ (11)
where $I_{0}(x)$ is the modified Bessel function of the first kind specialF . As is seen in Fig. 5, the approximation (9) accurately reproduces the exact results for the friction coefficient $\gamma>0.3$ which could be treated as the regime of strong damping. For the value $\gamma=1.66$ it is a local minimum of $D$, at this point $-J_{0}(\psi(1.66))\approx 0.4$ and the effective potential barrier assumes the maximal value. In turn, for $\gamma=3.3$ the diffusion coefficient is locally maximal and $J(\psi(3.3))\approx 0$, i.e. the effective potential barrier dissapears. It turns out that for larger values of the amplitude $a$ new maxima emerge. For example, if $a=18$ there are two maxima at $\gamma_{M}=1.05$ and $\gamma_{M}=1.4$ but for $a=35$ four maxima exist at $\gamma_{M}=0.95,2,3.75,9.5$ (not shown). The envelope of these points behaves roughly like $D(\gamma_{M})\sim 1/\gamma_{M}$.
For the case of weak dissipation, we can apply Eqs. (26) and (28) and then pavli ; entropy23
@@
where $d_{0}=J_{0}(\psi)$ and $\mathbf{E}(k)$ is the complete elliptic integral of the second kind specialF . In Fig. 5 we compare the dependence $D(\gamma)$ from this relation with the exact results of simulations. In this regime the discrepancy is very large and (12) cannot reconstruct the local maximum for $a=14$ which occurs at $\gamma=0.06$. Therefore this approximation totally fails. It is due to the fact that the periodic part of the ansatz (8) takes into account only locked trajectories but not the running ones. In the regime of strong damping, solely locked solutions exist (see panel (d) in Fig. 5) and then the approximation (8) is fine. On the other hand, in the regime of weak damping, there are also running trajectories like in Eq. (5), cf. panel (b) in Fig. 5. Therefore the ansatz (8) fails and in consequence Eq. (9) does not reproduce properties of the starting system (3). Its correct formulation is an interesting problem but lying beyond the scope of the current paper.
Discussion
@@
We showed that the dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ characterizing the dissipation mechanism can render unexpected features. Typically, $D$ is a decreasing function of $\gamma$. In contrast, we demonstrated that for an archetypal model of a nonequilibrium system consisting of a Brownian particle moving in a symmetric and spatially periodic potential subjected to an unbiased time-periodic driving the diffusion coefficient $D$ can be amplified up to several orders if the friction coefficient $\gamma$ is increased by a little bit. It can be realized for sufficiently large amplitude of the ac-driving and in the strong damping regime.
We revealed the mechanisms standing behind this counterintuitive phenomenon which is related to bifurcation of periodic orbits oscillating symmetrically around the potential maximum and their displacement towards the potential minimum when the friction coefficient increases. It induces the change of the effective barrier of the periodic potential and its vanishing at the maximum of $D$. Simultaneously, when $\gamma$ grows the intensity of thermal noise increases and, in turn, it intensifies activation processes and transitions between different states of the system. However, if damping is sufficiently strong, only decreasing of $D$ is observed as in the Einstein relation (1).
Properties in the strong dissipation regime are very well reconstructed by the vibrational mechanics methods proposed in Refs. boroEPL ; march-prl . On the other hand, in the weak dissipation regime, where the increase of diffusion vs friction is also observed, the effect is generated by a non-monotonic change of population of the running orbits. However, in this regime the enhancement of diffusion is small. Moreover, the vibrational mechanics approximation in the form (8)-(9) ceases to function in this regime.
We discussed only a single set of parameters $\{a,\omega,Q\}$ for which the giant acceleration of the diffusion process is detected. Are these three values exceptional? To answer this question we have searched the part of the parameter space in order to locate regions of giant enhancement of $D$. The inspection of the results revealed that there are many sets of parameter values for which this behavior occurs.
The phenomenon of giant enhancement of diffusion has been studied both theoretically and experimentally in a number of systems. In Ref. peter-epl the overdamped Brownian particle moving in a piecewise linear spatially periodic potential and driven by a piecewise constant time periodic driving was considered. It was shown that the diffusion coefficient exhibits quasi-periodic behavior as a function of the ”tilting time” of the piecewise constant force and at the maxima $D/D_{E}$ can be much larger than 1. In Refs. reimann2001a ; reimann2001b the overdamped dynamics of the Brownian particle in the washboard potential was studied and it was shown analytically that in a critically biased periodic potential the diffusion coefficient as a function of a constant force may be, in principle, arbitrarily enhanced. For a realistic experimental setup, an enhancement by 14 orders of magnitude was predicted so that thermal diffusion could be observable on a macroscopic scale at room temperature. The extension to the underdamped regime has been presented in Refs. marchenko2012 ; lindner16 and the regions of parameters for which the diffusion coefficient exponentially grows with inverse temperature have been identified. In Ref. urbakh it was demonstrated that lateral vibrations of the substrate can dramatically increase surface diffusivity and reduce friction at the nanoscale. It was analyzed in the context of atomic force microscopy by using the model of a tip interacting with a substrate, which oscillates in the lateral direction. The motion of the tip in the lateral, and normal, directions is governed by the coupled Langevin equations and it was found a sharp transition from the state with a small tip-surface separation to the state with a large separation as the vibration frequency increases.
The giant diffusion has been confirmed by experiments. In Ref. grier , a single colloidal sphere circulating around a periodically modulated optical vortex trap exhibits normal diffusion of the Eisnstein-like law with an effective diffusion coefficient that is enhanced by more than 2 orders of magnitude. In the next experiment nano , the diffusion of a gold nanoparticle in the nonconservative force field of an optical vortex lattice was observed and radiation pressure in the vortex array is shown to induce an enhancement by 2 orders of magnitude over the free thermal diffusion. In Ref. hayashi , diffusion to a single-molecule experiment on a rotary motor protein, $F_{1}$-ATPase , which is a component of $F_{0}F_{1}$ adenosine triphosphate synthase, was investigating. The rotation of $F_{1}$ was induced by applying an external torque using a single-molecule technique. It was found that the diffusion coefficient of a probe attached to $F_{1}$ (as a function of the external torque) shows a resonance peak, which was predicted by the theory of the giant acceleration of diffusion in Refs. reimann2001a ; reimann2001b . In the recent experiment rotor , the setup consists of a vacuum optical tweezer trapping a dielectric silica nanodumbbell. Its motion is driven by an elliptically polarized light beam tilting the angular potential. By varying the gas pressure (i.e. the friction) around the point of maximum intermittency, the effective diffusion coefficient increases by more than 3 orders of magnitude over free-space diffusion.
Properties of the studied system could be corroborated experimentally in several setups. We would like to mention two of them. The first is a resistively and capacitively shunted Josephson junction device. Dynamics of the phase difference $\Psi=\Psi(t)$ between the macroscopic wave functions across the Josephson junction in the framework of the Stewart-McCumber model stewart ; mccumber ; kautz1996 is described by the following equation
$\Big{(}\frac{\hbar}{2e}\Big{)}^{2}C\>\ddot{\Psi}+\Big{(}\frac{\hbar}{2e}\Big{)}^{2}\frac{1}{R}\dot{\Psi}+\frac{\hbar}{2e}I_{0}\sin\Psi=\frac{\hbar}{2e}I(t)+\frac{\hbar}{2e}\sqrt{\frac{2k_{B}T}{R}}\>\xi(t).$ (13)
The left hand side contains three additive contributions: a displacement current due to the capacitance $C$ of the junction, a normal (Ohmic) current characterized by the normal state resistance $R$ and a Cooper pair tunnel current characterized by the critical value $I_{0}$. In the right hand side, $I(t)$ is an external current applied to the device. There is an evident correspondence between the dimensional Langevin equation (14) (see below) and Eq. (13): the coordinate $x=\Psi$, the mass $M=(\hbar/2e)^{2}C$, the friction coefficient $\Gamma=(\hbar/2e)^{2}(1/R)$, the barrier height $\Delta U=(\hbar/2e)I_{0}$ and the period $L=2\pi$. Note that the increase of the friction $\Gamma$ now means the decrease of the resistance $R$. The particle velocity $v=\dot{x}$ corresponds to the voltage drop $V$ across the junction and the generalized Green-Kubo relation for the periodically driven processes allows to investigate the diffusion coefficient, see Appendix in Ref. jpc . Moreover, in this context it is worth mentioning the recent experiments concerning measurements of phase dynamics in Josephson Junctions and SQUIDs phase which open another possibility to test our findings.
The second example is related to an experimental setup based on the optical trapping of a nanodumbbell in a moderate vacuum. The system is trapped in an elliptically polarized laser beam rotor . For linearly polarized optical tweezers the dumbbell is subjected to a periodic force $\sin(x)$, cf. Eq. (1) in Ref. rotor . A special feature of this experiment is that the friction coefficient being linearly proportional to the gas pressure can be tuned over several orders of magnitude. It makes this optical trapping setup a suitable platform to study stochastic dynamics in the low-friction and strong-friction regimes. Recently, an effective technique called the optical feedback trap has been proposed feedback . It allows to create, with a high precision position detection and ultrafast feedback control, a mathematically driven spatio-temporal effective potential of any desired shape using the optical feedback force.
Our results could be applied, for instance, in microfluidic devices for controlled separation of particles through diffusion separator . The main idea is the following. There exist two species of particles with the linear sizes $R_{1}$ and $R_{2}=2R_{1}$. From the Stokes relation, $\gamma\propto R$, the corresponding friction coefficients for them are $\gamma_{1}$ and $\gamma_{2}=2\gamma_{1}$, and the diffusion coefficients are $D_{1}$ and $D_{2}$, respectively. If $\gamma_{1}\approx 1.65$ (see Fig. 2) then $\gamma_{2}\approx 3.3$ and $D_{2}\approx 5.6\times 10^{4}\,D_{1}$. From the relation (4) it follows that, statistically, if the smaller particles overcome the distance $X_{1}$ then the larger ones overcome the distance $X_{2}\approx\sqrt{5.6\times 10^{4}}\,X_{1}=2.4\times 10^{2}\,X_{1}$, i.e. two orders of magnitude greater. If the separator is properly designed, the larger particles can be effectively separated from the smaller ones. Recent developments of size-selective separation methods for nanoparticles using the external fields, filtration, and the stability of the colloid system are reviewed in Ref. mori .
Methods
Dimensional Langevin equation and scaling
@@ Methods
The dimensionless Langevin equation (3) is obtained from the dimensional Langevin equation in the form entropy23 ; marchPRE23 ; spiechowicz2016njp
@@
$F(t)=A\sin{(\Omega t+\phi_{0})},$ (16)
where $\phi_{0}$ is the initial phase. Thermal equilibrium fluctuations describing the influence of thermostat of temperature $T$ are normally assumed as $\delta$-correlated Gaussian white noise of zero-mean value,
@@
The characteristic length $L$ is the period of the potential $U(x)$ and the characteristic time $\tau_{0}$ is related to the period of small oscillations inside the well of $U(x)$.
@@
$\ddot{\hat{x}}+\gamma\dot{\hat{x}}=-\sin{\hat{x}}+a\sin(\omega\hat{t}+\phi_{0})+\sqrt{2\gamma Q}\,\hat{\xi}(\hat{t}),$ (19)
@@
$\gamma=\frac{\Gamma L}{2\pi\sqrt{M\Delta U}},\quad a=\frac{1}{2\pi}\frac{L}{\Delta U}A,\quad\omega=\Omega\frac{L}{2\pi}\sqrt{\frac{M}{\Delta U}},\quad Q=\frac{k_{B}T}{\Delta U}.$ (20)
The dimensionless potential $W(\hat{x})$ of the period $L=2\pi$ is $W(\hat{x})=U((L/2\pi)\hat{x})/\Delta U=-\cos{\hat{x}}$ and the corresponding potential force reads $f(\hat{x})=-W^{\prime}(\hat{x})=-\sin\hat{x}$. The rescaled thermal noise is $\hat{\xi}(\hat{t})=(L/2\pi\Delta U)\xi(t)=(L/2\pi\Delta U)\xi(\tau_{0}\hat{t})$ and has the same statistical properties as $\xi(t)$; i.e., $\langle\hat{\xi}(\hat{t})\rangle=0$ and $\langle\hat{\xi}(\hat{t})\hat{\xi}(\hat{s})\rangle=\delta(\hat{t}-\hat{s})$. The dimensionless noise intensity $Q$ is the ratio of thermal energy and half of the activation energy the particle needs to overcome the nonrescaled potential barrier. In order to simplify the notation we omit the hat-notation in Eq. (3).
The main problem studied in the paper is the dependence of the dimensionless diffusion coefficient $D$ on the dimensionless friction coefficient $\gamma$, see Eq. (20). If we want to consider a regime of the strong damping $\gamma\gg 1$ then it can be realized in two ways: (i) for fixed $M$ the friction coeffcient $\Gamma$ is large or (ii) for fixed $\Gamma$ the mass $M$ is small. If the control parameter is the dimensional friction coefficient $\Gamma$ then from Eq. (20) it follows that it is equivalent to the change of the dimensionless $\gamma$ while other parameters are fixed. However, if the control parameter is the dimensional mass $M$ then from Eq. (20) it follows that both $\gamma$ and $\omega$ are changed simultaneously. Therefore these two scenarios leading to the regime of strong damping are not equivalent if the remaining parameters are supposed to be fixed.
Diffusion coefficient: Selected analytical results
An analytical expression for the diffusion coefficient $D$ is here presented for some selected models. For the force-free Brownian particle or under the influence of a constant force $F$, i.e. when Eq.(3) assumes the form
$\ddot{x}+\gamma\dot{x}=F+\sqrt{2\gamma Q}\,\xi(t),$ (21)
the diffusion coefficient it given by the Einstein relation entropy23
$D_{E}=\frac{Q}{\gamma}.$ (22)
If the Brownian particle moves in random medium and this problem is modeled in terms of the fluctuating viscous friction
$\gamma\to\gamma(t)=\gamma_{0}+z(t)>0,$ (23)
where fluctuations $z(t)$ are assumed to be the zero-mean stationary random process $\langle z(t)\rangle=0$, then the diffusion coefficient is given by the expression Talkner1 ; Talkner2
[…]
where averaging on the right-hand side has to be performed over all realizations of fluctuations $z(t)$. Applying the Jensen inequality $\langle\exp(\xi)\rangle\geq\exp(\langle\xi\rangle)$, we note that $\tilde{D}$ is always greater than the Einstein diffusion coefficient.
$\ddot{x}+\gamma\dot{x}=-d_{0}\sin{x}+\sqrt{2\gamma Q}\,\xi(t),$ (25)
$D=\frac{Q}{\gamma}\,B(Q).$ (26)
In the case of overdamped dynamics ($\gamma\gg 1$), i.e. when the inertial term in Eq. (25) is formally neglected, it assumes the following form lifson ; entropy23
$B(Q)=\frac{1}{I_{0}^{2}(d_{0}/Q)}.$ (27)
$B(Q)=\frac{\sqrt{\pi d_{0}/Q}}{2I_{0}(d_{0}/Q)}\int_{1}^{\infty}\frac{\mbox{e}^{-d_{0}y/Q}\,dy}{\sqrt{y+1}\,\mathbf{E}(\sqrt{2/(y+1)})},$ (28)
where $I_{0}(x)$ is the modified Bessel function of the first kind and $\mathbf{E}(k)$ is the complete elliptic integral of the second kind specialF . In both cases $B(Q)$ depends only on $d_{0}$ and the dimensionless temperature $Q$. If $d_{0}$ does not depend on other parameters of the system then the diffusion coefficient (26) is smaller than the Einstein diffusion coefficient (22), i.e. $D\leq D_{E}$. Eq. (25) has the same form as Eq. (9) in which $d_{0}=J_{0}(\psi)$ and $\psi=\psi(a,\omega,\gamma)$. Therefore dependence of $D$ on $\gamma$ is more complicated than in Eq. (22). We stress that Eq. (25) is applied for verification of approximations in the strong and weak dissipation regimes, see subsection: Vibrational mechanics approach.
Data availability
The datasets generated during the current study are available from the corresponding author on reasonable request.
Acknowledgments
Author contributions
Conceptualization, discussion, writing and editing of the manuscript: I.G.M., J.Ł. and J.S.; simulations and calculations: I.G.M, V.A., I.I.M. and J.S. All authors have read and agreed to the published version of the manuscript.
Competing interests
The authors declare no competing interest
Additional information
Correspondence and requests for materials should be addressed to J.Ł. (e-mail: jerzy.luczka@us.edu.pl)